# VGBench: Evaluating Large Language Models on Vector Graphics Understanding and Generation

### v1

@@ Abstract
In the realm of vision models, the primary mode of representation is using pixels to rasterize the visual world. Yet this is not always the best or unique way to represent visual content, especially for designers and artists who depict the world using geometry primitives such as polygons. Vector graphics (VG), on the other hand, offer a textual representation of visual content, which can be more concise and powerful for content like cartoons or sketches. Recent studies have shown promising results on processing vector graphics with capable Large Language Models (LLMs). However, such works focus solely on qualitative results, understanding, or a specific type of vector graphics. We propose VGBench, a comprehensive benchmark for LLMs on handling vector graphics through diverse aspects, including (a) both visual understanding and generation, (b) evaluation of various vector graphics formats, (c) diverse question types, (d) wide range of prompting techniques, (e) under multiple LLMs. Evaluating on our collected 4279 understanding and 5845 generation samples, we find that LLMs show strong capability on both aspects while exhibiting less desirable performance on low-level formats (SVG). Both data and evaluation pipeline will be open-sourced at https://vgbench.github.io.
@@ 1 Introduction
Current vision models are mostly built on pixels, rasterizing the visual world into a matrix representation. Such rasterziation represents diverse visual content with equally sized elements. But pixels are not the only way to represent the visual world. For simple content such as cartoons or sketches, a different representation using explicit geometry primitives can be more concise and beneficial. Vector graphics offer such a textual representation for visual content via geometry primitives, e.g., circles and polygons, as shown in Figure 1 (a). Vector graphics have been critical for designers and artists since the geometry primitives can be easily manipulated. Vector representations include Scalable Vector Graphics (SVG), TikZ, Graphviz, etc.
@@
In this paper, we propose VGBench to comprehensively evaluate LLMs’ vector graphics processing capabilities via different aspects: VGBench (1) includes both visual understanding (VGQA) and generation (VGen); (2) evaluates diverse vector graphics formats such as SVG, TikZ, and Graphviz; (3) covers a set of taxonomies from low-level vision to high-level semantics, from color, shape, to category and advanced reasoning questions such as usage and the relation between objects; (4) adopts a variety of prompting techniques, such as zero-shot prediction, chain-of-thought reasoning, in-context learning, etc.; and (5) evaluates diverse LLMs including GPT-4 OpenAI (2023b), GPT-3.5 OpenAI (2023a), Llama-3-8B-Instruct, and Llama-3-70B-Instruct Meta (2024).
@@ 2.2 Evaluation for Image Understanding and Generation
Works on Image Understanding are mainly based on raster images. VQA Antol et al. (2015) first introduced the task of free-form and open-ended Visual Question Answering and evalauted existing LSTM-CNN based methods. CLIP Radford et al. (2021) introduces two encoders for both texts and images to achieve an aligned representation to serve as a baseline for many image understanding tasks. LLaVA Liu et al. (2024) and LLaMA-Adapter Zhang et al. (2023) propose approaches to solve general-purpose visual and language understanding problems based on large language models.
@@ 3.2 Experiment Settings
We primarily use GPT-4 (1106 version) OpenAI (2023b) as the medium for vector graphics understanding and generation. This is because GPT-4 shows superior language reasoning and generation capabilities, as previously mentioned in Section 2.3. We also evaluate GPT-3.5 OpenAI (2023a) and other open-source LLMs that are highly capable, including Llama-3-70B-Instruct Meta (2024) and Llama-3-8B-Instruct.
@@ 3.3 VGQA: Vector Graphics Understanding Benchmark
Finally, we collect 4279 samples in total, as shown in Table 1 and 2. The word distribution of answers in the VGQA dataset is illustrated in Figure 4. Specifically, we have 2228, 1139, and 912 samples for SVG, TikZ, and Graphviz, respectively. After an LLM makes responses to the vector graphics questions, we compare the final responses with the ground-truth answers to compute accuracy.
@@ 3.4 VGen: Vector Graphics Generation Benchmark
Llama-3-8B 32.3 39.8 48.0 40.0 64.6 53.0 45.9 54.5 68.0 52.5 55.8 58.8
Table 7: Evaluation of VGQA across diverse vector graphics formats for different LLMs in the zero-shot setting.
@@ 4 Conclusion
Our study unveils new insights into the capabilities of LLMs in understanding and generating vector graphics. We discovered that LLMs demonstrate decent vector graphics understanding in TikZ, Graphviz, and SVGs, with a particular strength in understanding vector graphics code with higher-level semantics. We also found that LLMs often exhibit strong vector graphics generation capabilities. Interestingly, advanced prompting techniques can significantly improve performance for low-level formats such as SVG, and while GPT-4 had the strongest performance, open-source models like Llama-3-70B show competitive performance. Our work lays a groundwork for future studies into LLMs’ vector graphics understanding and generation benchmarking, and we hope it will inspire further efforts to enhance these capabilities. We will release our benchmark dataset and evaluation pipeline.

### v2

@@ Abstract
In the realm of vision models, the primary mode of representation is using pixels to rasterize the visual world. Yet this is not always the best or unique way to represent visual content, especially for designers and artists who depict the world using geometry primitives such as polygons. Vector graphics (VG), on the other hand, offer a textual representation of visual content, which can be more concise and powerful for content like cartoons, sketches and scientific figures. Recent studies have shown promising results on processing vector graphics with capable Large Language Models (LLMs). However, such works focus solely on qualitative results, understanding, or a specific type of vector graphics. We propose VGBench, a comprehensive benchmark for LLMs on handling vector graphics through diverse aspects, including (a) both visual understanding and generation, (b) evaluation of various vector graphics formats, (c) diverse question types, (d) wide range of prompting techniques, (e) under multiple LLMs and (f) comparison with VLMs on rasterized representations. Evaluating on our collected 4279 understanding and 5845 generation samples, we find that LLMs show strong capability on both aspects while exhibiting less desirable performance on low-level formats (SVG). Both data and evaluation pipeline will be open-sourced at https://vgbench.github.io.
@@ 1 Introduction
Current vision models are mostly built on pixels, rasterizing the visual world into a matrix representation. Such rasterized represents diverse visual content with equally sized elements. But pixels are not the only way to represent the visual world. For contents such as cartoons, sketches or scientific figures, a different representation using explicit geometry primitives can be more concise and beneficial. Vector graphics offer such a textual representation for visual content via geometry primitives, e.g., circles and polygons, as shown in Figure 1 (a). Vector graphics have been critical for designers and artists since the geometry primitives can be easily manipulated. Vector representations include Scalable Vector Graphics (SVG), TikZ, Graphviz, etc.
@@
In this paper, we propose VGBench to comprehensively evaluate LLMs’ vector graphics processing capabilities via different aspects: VGBench (1) includes both visual understanding (VGQA) and generation (VGen); (2) evaluates diverse vector graphics formats such as SVG, TikZ, and Graphviz; (3) covers a set of taxonomies from low-level vision to high-level semantics, from color, shape, to category and advanced reasoning questions such as usage and the relation between objects; (4) adopts a variety of prompting techniques, such as zero-shot prediction, chain-of-thought reasoning, in-context learning, etc.; (5) evaluates diverse LLMs including GPT-4 OpenAI (2023b), GPT-3.5 OpenAI (2023a), Llama-3-8B-Instruct, Llama-3-70B-Instruct Meta (2024), Qwen2-7B-Instruct, Qwen2-72B-Instruct qwe (2024), Phi-3-mini-128k-instruct, Phi-3-medium-128k-instruct Abdin et al. (2024), gemini-1.5-pro Reid et al. (2024); and (6) evaluates the VLM LLaVA-1.5-13b Liu et al. (2024a) over rasterized representations of images in our benchmark..
@@ 2.2 Evaluation for Image Understanding and Generation
Works on Image Understanding are mainly based on raster images. VQA Antol et al. (2015) first introduced the task of free-form and open-ended Visual Question Answering and evalauted existing LSTM-CNN based methods. CLIP Radford et al. (2021) introduces two encoders for both texts and images to achieve an aligned representation to serve as a baseline for many image understanding tasks. LLaVA Liu et al. (2024b) and LLaMA-Adapter Zhang et al. (2023) propose approaches to solve general-purpose visual and language understanding problems based on large language models.
@@ 3.2 Experiment Settings
We primarily use GPT-4 (1106 version) OpenAI (2023b) as the medium for vector graphics understanding and generation. This is because GPT-4 shows superior language reasoning and generation capabilities, as previously mentioned in Section 2.3. We also evaluated other proprietary models such as GPT-4o, GPT-3.5 OpenAI (2023a) and gemini-1.5-pro Reid et al. (2024), along with many other open-source LLMs that are highly capable, including Llama-3-70B-Instruct Meta (2024), Llama-3-8B-Instruct, Qwen2-7B-Instruct, Qwen2-72B-Instruct qwe (2024), Phi-3-mini-128k-instruct and Phi-3-medium-128k-instruct Abdin et al. (2024).
@@ 3.3 VGQA: Vector Graphics Understanding Benchmark
Finally, we collect 4279 samples in total, as shown in Table 1 and 2. The word distribution of answers in the VGQA dataset is illustrated in Figure 4. Specifically, we have 2228, 1139, and 912 samples for SVG, TikZ, and Graphviz, respectively. After an LLM makes responses to the vector graphics questions, we compare the final responses with the ground-truth answers to compute accuracy. For LLMs with weaker instruction-following capabilities in producing easily parsable outputs, we use GPT-4 to determine their chosen option and then assess their accuracy.
@@ 3.4 VGen: Vector Graphics Generation Benchmark
Qwen2-7B 33.3 48.7 46.3 42.8 79.4 64.7 58.3 67.5 81.8 57.3 68.6 69.2
Qwen2-72B 43.4 62.4 55.9 53.9 88.6 74.6 72.5 78.6 86.5 71.5 80.8 79.6
Phi-3-Mini-128K 34.1 29.8 49.7 37.9 70.6 52.5 50.7 57.9 74.7 58.9 68.6 67.4
Phi-3-Medium-128k 43.6 44.7 60.6 49.6 80.4 59.7 62.8 67.6 81.4 66.5 72.7 73.5
Gemini-1.5-Pro 39.2 73.2 47.9 53.4 86.7 74.9 71.8 77.8 79.5 66.8 86.0 77.4
LLaVA-1.5-13b-hf 83.0 85.2 84.0 84.1 64.3 34.3 44.8 47.8 46.7 53.9 49.7 50.1
Table 7: The evaluation of VGQA across diverse vector graphics formats for different LLMs and the evaluation of rasterized representation of VGQA in VLMs in the zero-shot setting.
@@ 3.5 In Depth Analysis
Comparison between LLMs and MLLMs on Image Understanding
We find that VLMs such as LLaVA Liu et al. (2023) show interesting behavior compared with LLMs on vector graphics. To evaluate the performance of VLMs on VGQA, we render each visual content in our benchmark into the PNG format, and then feed the same question to VLMs. Specifically, we evaluate LLaVA-1.5-13bLiu et al. (2024a), as shown in Table 9. LLaVA-1.5 shows strong performance in SVG format rather than TikZ and Graphviz. The strong performance gain that LLMs obtained on high-level vector graphics languages such as TikZ and Graphviz shows that those kinds of vector graphics are more aligned with LLMs’ training data, natural languages, which is a highly compressed representation of the world. Low-level vector graphics languages such as SVG cover more low-level visual signals that can be better handled by VLMs using their rasterized representation.
LLM SVG Avg TikZ Avg Graphviz Avg
GPT-4 54.9 81.0 84.5
GPT-4o(Text) 64.4 79.8 80.8
LLaVA-1.5-13b 84.1 47.8 50.1
Table 9: The comparison on image understanding ability between LLMs and VLMs, reflected by the average accuracy (%) on VGQA. We feed the rasterialized vector graphics in PNG format along with the same question in text format to LLaVA-1.5-13b to evaluate its performance on VGQA
@@ 4 Conclusion
Our study unveils new insights into the capabilities of LLMs in understanding and generating vector graphics. We discovered that LLMs demonstrate decent vector graphics understanding in TikZ, Graphviz, and SVGs, with a particular strength in understanding vector graphics code with higher-level semantics. We also found that LLMs often exhibit strong vector graphics generation capabilities. Interestingly, advanced prompting techniques can significantly improve performance for low-level formats such as SVG, and while GPT-4 had the strongest performance, open-source models like Llama-3-70B and Qwen2-72B show competitive performance. Our work lays a groundwork for future studies into LLMs’ vector graphics understanding and generation benchmarking, and we hope it will inspire further efforts to enhance these capabilities. We will release our benchmark dataset and evaluation pipeline.